OpenStudy (anonymous):

How do you write a equation of slope -5 passes through (2,-3)

OpenStudy (anonymous):

Substitute the values into the equation \(y = mx + b\): $$-3 = -5(2) + b$$ Solve for b, and then rewrite just in general.

OpenStudy (anonymous):

Okay and then how would you for passes through (-1,8) (5,4)

OpenStudy (anonymous):

Alright, so we're not given the slope. However, slope is defined by the rate of change between any two points: $$m = \frac{-1 - 5}{8- 4}$$ Now, given the slope, repeat the process we did for the first equation.

OpenStudy (anonymous):

Okay thank you